Assume that today is December 31, 2019, and that the following information applies to Abner Airlines: After-tax operating income [EBIT(1 - T)] for 2020 is expected to be $700 million. The depreciation expense for 2020 is expected to be $120 million. The capital expenditures for 2020 are expected to be $475 million. No change is expected in net operating working capital. The free cash flow is expected to grow at a constant rate of 6% per year. The required return on equity is 13%. The WACC is 9%. The firm has $202 million of non-operating assets. The market value of the company's debt is $3.704 billion. 160 million shares of stock are outstanding. Using the corporate valuation model approach, what should be the company's stock price today? Do not round intermediate calculations. Round your answer to the nearest cent.
Answer #1

Free Cash Flow (FCF) = After-tax operating income + depreciation - capex

= 700 + 120 -475 = 345 million

Value of operations (Vo) = FCF/(WACC -g) = 345/(9%-6%) = 11,500 million

Firm value = value of operations + non-operating assets

= 11,500 + 202 = 11,702 million

Equity value = Firm value - debt = 11,702 - 3,704 = 7,998 million

Price per share = equity value/number of shares outstanding = 7,998/160 = $49.99